A particle with charge 5.0-cu is placed at the corner of a cube. The total electric flux ON maC through all sides of the cube is


1
Expert's answer
2020-11-25T07:13:47-0500

Flux through any surface is given by;

"\Phi = \frac{Q}{\epsilon_{o}}"

If a corner atom is shared by eight neighboring such cubes, then one-eighth part of the corner atom or particle contributes to each cube.

"\Phi = \frac{Q}{8\epsilon_{o}}=\frac{5\times 10^{-6}}{8\times 8.85 \times 10^{-12}} = 7.1 \times 10^{4}Nm^{2}/C"